OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

what have u tried so far?

OpenStudy (anonymous):

|x − 3| = 6 x − 3 = ± 6 x = 3 ± 6

OpenStudy (anonymous):

Am really confused.

OpenStudy (anonymous):

Are you sure you typed that right? There aren't any solutions for x which satisfy that equation

OpenStudy (anonymous):

Sorry, I made a mistake :( x can be any real number because if you multiply out the LHS you get: -2x -6 and if you do the same for the RHS you get: -2x -6, therefore, -2x-6=-2x-6 which is always true

OpenStudy (anonymous):

okay so then I work out both sides.

OpenStudy (anonymous):

multiply out the brackets on both sides

OpenStudy (anonymous):

You get it?

OpenStudy (anonymous):

1. -2(x+3) = -2(x+1) - 4 DISTRIBUTE 2. -2x - 6 = -2x - 2 - 4 COMBINE LIKE TERMS 3. -2x - 6 = -2x - 6 As you can see, these are the same equations. X = all real numbers

OpenStudy (anonymous):

As you can see karen , the LHS balances with the RHS as Cupcakesforall6 has shown. hence x can be solution of every element from the set of real numbers. {x:x x∈R}